### v1

# CoT-based Synthesizer: Enhancing LLM Performance through Answer Synthesis

Abstract
Current inference scaling methods, such as Self-consistency and Best-of-N, have proven effective in improving the accuracy of LLMs on complex reasoning tasks. However, these methods rely heavily on the quality of candidate responses and are unable to produce correct answers when all candidates are incorrect. In this paper, we propose a novel inference scaling strategy, CoT-based Synthesizer, which leverages CoT reasoning to synthesize superior answers by analyzing complementary information from multiple candidate responses, even when all candidate responses are flawed. To enable a lightweight and cost-effective implementation, we introduce an automated data generation pipeline that creates diverse training data. This allows smaller LLMs trained on this data to improve the inference accuracy of larger models, including API-based LLMs. Experimental results across four benchmark datasets with seven policy models demonstrate that our method significantly enhances performance, with gains of 11.8% for Llama3-8B and 10.3% for GPT-4o on the MATH dataset. The corresponding training data and code are publicly available on the repository [^1] https://github.com/RUCKBReasoning/CoT-based-Synthesizer.
CoT-based Synthesizer: Enhancing
LLM Performance through Answer Synthesis
Bohan Zhang1,3thanks: Equal Contributions.thanks: Work was done when interned at Zhipu AI., Xiaokang Zhang1,311footnotemark: 1, Jing Zhang1,3thanks: Corresponding Author., Jifan Yu2, Sijia Luo1, Jie Tang2 1School of Information, Renmin University of China, 2Tsinghua University, 3Key Laboratory of Data Engineering and Knowledge Engineering, Beijing, China {zbhmint, zhang2718, zhang-jing}@ruc.edu.cn
1 Introduction
While large language models (LLMs) Brown et al. (2020); Chowdhery et al. (2023); Touvron et al. (2023); Achiam et al. (2023) have achieved revolutionary progress in the field of natural language processing (NLP), they often struggle to generate accurate answers in one attempt when dealing with complex reasoning tasks Rae et al. (2021); He et al. (2024); Tong et al. (2024). To address this, researchers have focused on expanding the scale of inference to enhance model inference performance Nye et al. (2021); Wu et al. (2024); Brown et al. (2024).
Figure 1: An example of our method in mathematical reasoning. Even when the policy model generates all incorrect responses, our method can still leverage their strengths to produce the correct answer.
A common inference scaling strategy is Best-of-N (BoN), which involves training a verifier to score multiple candidate responses generated by the model and selecting the highest-scoring answer as the final result Stiennon et al. (2020); Cobbe et al. (2021); Lightman et al. (2024). However, the scoring process for each candidate response in Best-of-N is independent, which fails to leverage the potential relationships between the candidates. This may lead to insufficient information integration and a susceptibility to reward hacking Skalse et al. (2022); Singhal et al. (2023). Since candidates are evaluated in isolation, the lack of cross-sample comparison makes it hard to detect outputs that exploit flaws in the reward function Christiano et al. (2017); Ouyang et al. (2022). Another widely used inference scaling strategy is Self-consistency (SC), which takes a distinct approach by selecting the most frequently generated answer from multiple inferences as the final result Wang et al. (2022). Unlike BoN, SC does not require an additional scoring step by a verifier. Instead, it uses a majority-voting process based on exact matches of candidate responses. However, this reliance on exact matching limits SC’s applicability in open-ended tasks. To address this limitation, the Universal Self-Consistency (USC) Chen et al. (2023) leverages the generative capabilities of LLMs for voting, replacing exact matching in SC, thus expanding its applicability across a wider range of tasks. However, when the correct answer appears at a low frequency among the candidates, consistency methods may fail to select the correct answer. In summary, both BoN and consistency methods operate on the assumption that the correct answer exists within the candidate set. When all candidate responses are incorrect, these methods fail to generate the correct answer. Given that BoN and USC both rely on post-processing models to evaluate candidate responses, a natural question emerges: Can we go beyond selecting from existing candidates and instead synthesize a better answer by combining the strengths of multiple candidate responses?
Based on this insight, we propose a novel inference scaling strategy: CoT-based Synthesizer, which leverages Chain-of-Thought (CoT) Wei et al. (2022) reasoning to systematically analyze candidate responses and synthesize new answers during the inference process. By doing so, the model can identify the strengths of each candidate and synthesize them into a more accurate and complete final answer. As illustrated in Figure 1, Best-of-N and Self-consistency fail to derive the correct answer from flawed candidates. In contrast, by encouraging the model to deeply analyze and integrate information across candidate responses, our method enables the synthesis of the correct answer. To enable training a smaller yet effective LLM for this purpose, we further propose an automated data generation pipeline designed to create training data for candidate response synthesis. By sampling multiple inference outputs from LLMs and filtering to ensure quality, we collect diverse data with complementary information for training.
We conduct experiments on 4 benchmarks and evaluate our method using 7 models with varying parameter scales for candidate response generation. The results demonstrate that small models trained using the data generation pipeline can not only effectively enhance the performance of candidate responses generated by various large models, even API models (e.g., with gains of $11.8\%$ for Llama3-8B and $10.3\%$ for GPT-4o on MATH), but also outperform traditional methods such as Self-consistency and Best-of-N.
Our contributions can be summarized as follows: 1) We propose a novel inference scaling strategy: CoT-based Synthesizer, which synthesizes better results by analyzing candidate responses. 2) We design an automated data generation pipeline to enable training smaller models for superior answer synthesis. 3) We validate the effectiveness of our method across multiple NLP tasks, achieving significant improvements over existing methods in complex reasoning scenarios.
2 Related Work
Best-of-N.
The Best-of-N approach, introduced as a common method for improving the quality of generated outputs, involves sampling multiple solutions and selecting the best one based on a scoring mechanism Stiennon et al. (2020). Previous works have trained a reward model, often referred to as a “verifier”, to discriminate between correct and incorrect solutions, either through binary classification Cobbe et al. (2021); Uesato et al. (2022); Yu et al. (2023) or by ranking preferences between solutions Stiennon et al. (2020); Nakano et al. (2021). Recent works integrate the generative capabilities of LLMs with the discriminative power of reward models, enabling both scoring and explaining the scoring process Zhang et al. (2024a); Yu et al. (2024). This paper focuses exclusively on outcome-based methods, leaving process-oriented approaches beyond the scope of discussion.
Consistency Methods.
Self-consistency is another common inference scaling method, which operates under the assumption that the most consistent answer, determined through a voting mechanism across multiple reasoning paths, is more likely to be correct. Methods based on consistency have demonstrated significant performance improvements across various domains like mathematics Wang et al. (2022), code generation Li et al. (2022); Shi et al. (2022) and open-ended question answering Chen et al. (2023). Our method shares similarities with Chen et al. (2023) in utilizing LLMs for inference scaling, while our method is distinguished by its emphasis on the generation of novel answers. Notably, even when all candidate responses are incorrect, our method retains the capability to yield accurate outcomes.
Inference Scaling Synthesis Methods.
Recent studies Farinhas et al. (2023); Vernikos et al. (2023) have explored answer synthesis by combining candidate responses. Additionally, Vernikos and Popescu-Belis (2024) introduces methods that leverage quality estimation metrics to efficiently aggregate outputs from LLMs. These approaches have successfully validated the efficacy of synthesizing high-quality responses in translation tasks. Similarly, ensemble methods have been explored for combining candidate responses from different LLMs. One such method trains a ranker to conduct pairwise comparisons and selects suitable candidates to fuse into a superior answer Jiang et al. (2023). However, our work focuses on synthesizing candidate responses generated exclusively by a single LLM without using diverse models.
3 Problem Formulation
This study concentrates on post-processing multiple candidate responses generated by the policy model to produce a new synthesized answer. In this paper, the policy model refers to the model that generates direct responses to user queries. Formally, given a user query $x$, the policy model performs $N$ rounds of random sampling-based generation from a probability distribution, producing a set of responses denoted as: $R=\{r_{1},r_{2},\dots,r_{N}\}$, where $r_{i}$ represents the $i$-th candidate answer. These responses, together with the query $x$, are then combined into $\{x,R\}$. Subsequently, an analysis and synthesis model, referred to as the Synthesizer, is employed to regenerate a new answer $y$. This process can be formalized as:
$(x,R)\xrightarrow{f_{\text{Synthesizer}}}y$
Here, $f_{\text{Synthesizer}}$ denotes the synthesis function applied to the query $x$ and candidate set $R$ to produce the final answer $y$.
4 Methodology
Figure 2: An overview of our method. (1) Synthesizer Inference: The policy model generates diverse candidate responses, which are analyzed and synthesized by the CoT-based synthesizer to produce a high-quality final response. (2) Data Generation Pipeline: The pipeline combines a diverse response generation process using a sampling LLM with query-response relationship analysis to construct high-quality synthetic data. (3) Synthesizer Training: The generated dataset is then used to train the model via SFT to enhance reasoning and synthesis capabilities.
As illustrated in Figure 2, we present an overview of our methodology. Specifically, we propose an inference scaling strategy based on the CoT reasoning, enabling the model to analyze and synthesize information across multiple candidate responses. To enable training a smaller yet effective LLM, we devise a two-stage data generation pipeline to construct a high-quality training dataset, which combines automated candidate generation with iterative repair mechanisms to improve data reliability. Finally, leveraging the generated dataset, we train Synthesizer-8B that integrates CoT reasoning to synthesize the final answer.
4.1 Synthesizer Inference
Diverse Response Generation.
We use the query as the input for the policy model to generate a set of candidate answers. To ensure diversity and quality among candidates, we employ the following decoding strategies during generation: (1) We set a relatively high sampling temperature ($t=0.9$) to increase randomness and promote diversity. (2) We employ Top-P ($p=0.9$) sampling decoding Holtzman et al. (2020), which truncates the candidate pool to include only the most probable tokens up to a cumulative probability of $p$. This process inherently avoids generating rare but semantically meaningless low-probability candidates, ensuring that the generated responses remain coherent and relevant. By combining these strategies, we generate a diverse yet coherent ensemble of $N$ responses, designated as $(x,R)$.
Response Analysis and Synthesis.
We input the generated $(x,R)$ into the CoT-based Synthesizer, which aims to analyze candidate responses, extract accurate information, and ultimately produce a high-quality response. The process involves the following steps: (1) Response Analyis: Synthesizer first conducts a thorough analysis of the relationship between the user query and each candidate response. It takes into account the frequency, relevance, and accuracy of similar responses to user queries. While high-frequency answers may indicate higher credibility, the Synthesizer does not overly rely on frequency as an evaluation criterion. Instead, it prioritizes logical coherence and factual accuracy, making it possible to identify even partially correct or misleading answers and extract valuable information from them. (2) Response Synthesis: When a correct answer exists within the candidate set, the Synthesizer further enhances and enriches the final response by incorporating effective reasoning steps from other candidate answers. In cases where there is a lack of clearly correct options or all candidate answers have deficiencies, the Synthesizer leverages its reasoning capabilities to integrate reasonable elements from multiple candidates, constructing a more coherent and accurate answer. These processes are achieved by designing specific prompts for the Synthesizer, enabling structured reasoning and accurate response generation. Detailed information can be seen in Box C.
4.2 Data Generation Pipeline
To enable the training of CoT-based Synthesizer, we design a two-stage data generation pipeline. Our pipeline is built on top of existing benchmarks, where we denote the query in the benchmark as $x$ and its corresponding gold answer as $y_{g}$.
Synthetic Answer Generation.
For a given training query $x$, we follow the diverse response generation process described in Section 4.1, using a sampling LLM (e.g., Llama3-8B-Instruct) to obtain the pair $(x,R)$. Then we input the pair into a high-performance response LLM (e.g., Llama3.1-70B-Instruct). Given the complexity of $x$, the response LLM may not synthesize an accurate answer on the first attempt. To address this problem, we perform multiple samplings, significantly increasing the likelihood of obtaining a correct response Tong et al. (2024). Specifically, we sample the response LLM $N$(e.g., $N=50$) times to conduct response analysis and synthesis as described in Section 4.1, generating a diverse set of synthetic answers. To ensure the quality of these synthetic answers, we filter responses using the gold answer $y_{g}$, retaining correct responses $y$, which contain CoT analysis and synthesis of $R$.
LLM Repair.
In cases where all candidate responses $R$ are incorrect, it is challenging for the response model to generate a correctly synthesized answer during the initial sampling stage. Specifically, we explicitly inform the response model that all candidate responses $R$ are incorrect and prompt it to reflect on the errors in these responses. By analyzing the incorrect answers, the model identifies valid reasoning steps and synthesizes a refined response that is more likely to be correct. Then we filter the newly generated responses to obtain the correct responses $y$.
4.3 Synthesizer Training
Through the above stages, we construct the synthesis training dataset $M=\{(x^{(i)},R^{(i)},y^{(i)})\}_{i=1}^{N}$, where $x$ represents the user query, $R$ is the set of candidate responses, and $y$ is the correct synthesis answer generated by the Response LLM. Based on this dataset, we train the model CoT-based Synthesizer-8B. The model takes the user query $x$ and the candidate response set $R$ as input and generates the correct answer $y$ as output. The generation probability is defined as:
$p_{\phi}(y\mid x,R)=\prod_{i=1}^{T}p_{\phi}(y_{i}\mid x,R,y_{<i}),$ (1)
where $T$ is the sequence length, $y_{i}$ is the $i$-th target token, and $y_{<i}$ represents the generated context (i.e., all previous tokens).
Through training, the model learns to leverage both the user query $x$ and the candidate response set $R$ to generate accurate and contextually appropriate answers $y$. This process enhances the model’s ability to synthesize coherent outputs while effectively utilizing the provided candidates.
5 Experiment
Table 1: The main result of our approach and other baselines over GSM8k, MATH, WikiTQ and FeTaQA. The top two performances are highlighted in bold and underlined.
Method Policy Models
GLM-4-Plus GPT-4o Llama3.1-70B Llama3.1-8B Llama3-8B Qwen2-7B Qwen2.5-14B Average
GSM8k
CoT-prompting Wei et al. (2022) 88.6 91.4 92.7 81.9 73.4 82.0 91.2 85.9
SC Wang et al. (2022) 90.1 92.4 93.9 85.1 80.9 84.3 92.3 88.4
USC(Llama3.1-70B) Chen et al. (2023) 90.1 92.3 93.5 85.4 82.0 84.9 92.3 88.6
LMCOR(Llama3-8B) Vernikos et al. (2023) 88.9 90.4 90.1 83.1 79.4 84.8 89.5 86.6
ArmoRM(Llama3-8B) Wang et al. (2024) 90.3 91.6 93.3 85.5 82.4 86.1 92.1 88.8
Scalar RM(Llama3-8B) Cobbe et al. (2021) 89.1 91.9 93.3 85.6 81.6 85.8 91.4 88.4
ours(Llama3.1-70B) 91.2 92.6 93.6 86.9 83.5 88.3 92.3 89.8
ours(Synthesizer-8B) 91.4 93.0 94.0 86.1 81.3 86.4 92.7 89.3
MATH
CoT-prompting Wei et al. (2022) 54.8 62.5 66.6 46.5 24.2 57.3 74.4 55.2
SC Wang et al. (2022) 63.0 68.7 68.8 55.4 32.4 61.0 76.6 60.8
USC(Llama3.1-70B) Chen et al. (2023) 62.6 67.3 68.4 52.8 35.4 62.2 78.2 61.0
LMCOR(Llama3-8B) Vernikos et al. (2023) 52.4 61.2 57.6 44.8 33.6 51.6 64.0 52.2
ArmoRM(Llama3-8B) Wang et al. (2024) 60.6 67.5 69.4 52.6 32.8 60.2 77.2 60.0
Scalar RM(Llama3-8B) Cobbe et al. (2021) 61.4 65.9 66.8 52.8 34.2 59.4 77.6 59.7
ours(Llama3.1-70B) 64.2 75.5 69.6 52.8 38.8 63.6 79.0 63.4
ours(synthesizer-8B) 64.4 72.8 69.6 54.6 36.0 62.4 78.2 62.6
WikiTQ
CoT-prompting Wei et al. (2022) 90.1 89.9 86.7 72.4 71.7 63.8 77.9 78.9
USC(Llama3.1-70B) Chen et al. (2023) 91.6 91.8 88.3 79.6 76.3 69.2 81.5 82.6
LMCOR(Llama3-8B) Vernikos et al. (2023) 88.8 90.4 87.8 77.3 75.4 69.2 81.4 81.5
ArmoRM(Llama3-8B) Wang et al. (2024) 91.0 91.8 87.5 77.9 73.8 69.4 81.2 81.8
Scalar RM(Llama3-8B) Cobbe et al. (2021) 91.8 90.5 87.5 77.6 74.9 69.8 80.1 81.7
ours(Llama3.1-70B) 91.9 92.3 88.3 83.4 82.2 78.0 84.2 85.8
ours(Synthesizer-8B) 92.1 91.9 88.9 79.9 77.7 72.2 82.4 83.6
FeTaQA
CoT-prompting Wei et al. (2022) 86.4 86.3 85.6 82.6 82.2 73.5 82.7 82.8
USC(Llama3.1-70B) Chen et al. (2023) 87.1 87.0 86.1 84.3 83.9 77.5 84.1 84.3
LMCOR(Llama3-8B) Vernikos et al. (2023) 86.0 84.7 83.0 84.7 83.8 79.9 83.2 83.6
ArmoRM(Llama3-8B) Wang et al. (2024) 87.5 86.1 86.0 83.2 82.5 76.1 82.9 83.5
Scalar RM(Llama3-8B) Cobbe et al. (2021) 87.4 85.5 85.3 83.0 82.3 75.1 83.3 83.1
ours(Llama3.1-70B) 87.0 86.8 86.6 84.9 85.1 82.3 84.1 85.3
ours(Synthesizer-8B) 87.5 87.9 87.5 84.7 85.9 82.1 86.6 86.0
Table 2: The impact of the data generation pipeline and the process of CoT analysis of candidate responses. The improvements are calculated between methods and CoT-prompting.
Policy Models
Method GLM-4-Plus GPT-4o Llama3.1-70B Llama3.1-8B Llama3-8B Qwen2-7B Qwen2.5-14B Average
GSM8k
CoT-prompting 88.6 91.4 92.7 81.9 73.4 82.0 91.2 85.9
Synthesizer-8B(ours) 91.4 +2.8 93.0 +1.6 94.0 +1.3 86.1 +4.2 81.3 +7.9 86.4 +4.4 92.7 +1.5 89.3 +3.4
     w/o CoT training 90.0 +1.4 90.6 -0.8 91.6 -1.1 83.3 +1.4 78.4 +5.0 84.3 +2.3 90.4 -0.8 86.9 +1.0
     w/o training 87.0 -1.6 87.3 -4.1 89.4 -3.3 80.1 -1.8 75.4 +2.0 81.8 -0.2 87.5 -3.7 84.1 -1.8
MATH
CoT-prompting 54.8 62.5 66.6 46.5 24.2 57.3 74.4 55.2
Synthesizer-8B(ours) 64.4 +9.6 72.8 +10.3 69.6 +3.0 54.6 +8.1 36.0 +11.8 62.4 +5.1 78.2 +3.8 62.6 +7.4
     w/o CoT training 57.6 +2.8 67.1 +4.6 63.4 -3.2 49.6 +3.1 37.0 +12.8 58.0 +0.7 71.4 -3.0 57.7 +2.5
     w/o training 59.2 +4.4 67.7 +5.2 64.6 -2.0 50.1 +3.6 32.2 +8.0 58.0 +0.7 70.6 -3.8 57.5 +2.3
5.1 Experimental Setup
Datasets and Metrics.
In this paper, we evaluate the Synthesizer on the following tasks:
  • Mathematical reasoning benchmarks, including GSM8k Cobbe et al. (2021) and MATH500 Lightman et al. (2024). GSM8k is a widely used dataset for grade school math problems, while MATH500 is a subset extracted from the MATH dataset Hendrycks et al. (2021), containing a diverse variety of challenging high school competition questions.
  • Table question answering (TableQA) benchmarks, including WikiTQ Pasupat and Liang (2015) and FeTaQA Nan et al. (2022), which are two widely used TableQA benchmarks.
We focus on these tasks because both require substantial reasoning capabilities, and current LLMs still exhibit significant room for improvement in these domains. For the TableQA datasets, which require preprocessing to ensure format compatibility, we directly utilized the processed WikiTQ and FeTaQA test sets released by TableLLM Zhang et al. (2024b). Since our model is fine-tuned to adhere to a specific format, we extract answers that align with these format requirements from the LLM’s outputs to perform an exact match (EM) evaluation against the gold answers. Therefore, on the GSM8k, MATH, and WikiTQ datasets, we use EM to calculate accuracy. Since the gold answers in the FeTaQA dataset are in the form of complete sentences, unlike the short answer phrases in WikiTQ, we select Rouge-L as the evaluation metric to better assess the quality of generated textual answers.
Baselines and Implementation.
We evaluate our method on the following models: API-based models (GPT-4o Achiam et al. (2023), GLM-4-plus GLM et al. (2024)) and open-source models (Llama3-8B-Instruct Dubey et al. (2024), Llama3.1-(8B,70B)-Instruct Dubey et al. (2024), Qwen2-7B-Instruct Yang et al. (2024), Qwen2.5-14B-Instruct Yang et al. (2024)).
The baselines for comparison are as follows:
  • CoT-prompting Wei et al. (2022): We prompt the policy model to directly generate a single response without any post-processing techniques.
  • Consistency-based methods: Self-consistency (SC) Wang et al. (2022) and Universal Self-consistency (USC) Chen et al. (2023). We use SC only on GSM8k and MATH since their answers are easy to extract in complete form, allowing majority voting based on exact match. Additionally, we adopt USC across all benchmarks.
  • Best-of-N methods: We first employ the ArmoRM Wang et al. (2024) built on Llama3-8B, which is the top-ranked Llama3-8B-based model on RewardBench Lambert et al. (2024). To ensure a fairer comparison, we also train task-specific scalar reward model (Scalar RM) for MATH and WikiTQ datasets, which is the same as that used in our Synthesizer training, following the method in Cobbe et al. (2021).
  • Synthesis methods: LMCOR Vernikos et al. (2023), which directly uses the gold answers $y_{g}$ from the benchmarks as the target responses for training its synthesis model.
For our method, as described in Section 4.2, we utilize Llama3-8B-Instruct to generate candidate responses and Llama3.1-70B-Instruct to produce synthetic answers. Data generation is conducted on the training sets of MATH and WikiTQ, tailored to mathematical reasoning and TableQA. During the data filtering stage, for mathematical reasoning tasks, we filter out incorrect answers by performing an exact match between the synthesized answers and the gold answer. For TableQA, we follow TableLLM Zhang et al. (2024b), using CritiqueLLM Ke et al. (2023) to score the synthesized answers and retain only high-quality responses based on the scores, as the generated answers are often lengthy and challenging to precisely match. Subsequently, we train the CoT-based Synthesizer on Llama3-8B-Instruct Dubey et al. (2024), as it demonstrates robust general capabilities and has not yet saturated the current datasets.
Moreover, for all the aforementioned methods, we use Llama3.1-70B-Instruct as the backbone for tuning-free methods, and Llama3-8B-Instruct for training-based methods. Responses are generated by sampling k = 5 times from the policy model with a temperature of 0.9 and a Top-P value of 0.9. Further details regarding experimental settings and implementation are provided in Appendix B.
5.2 Main Results
Table 1 presents the experimental results of across four benchmarks. The results demonstrate that our method significantly improves the reasoning performance of various policy models. Below, we summarize the key findings:
Our method consistently outperforms baselines across benchmarks. Across the four benchmarks, Self-consistency, ArmoRM, Scalar RM and our method generally improve the CoT outputs of various policy models. We evaluate a total of seven policy models. 1) On the MATH and WikiTQ datasets, our method (including both Llama3.1-70B-Instruct and Synthesizer-8B) outperforms all baseline methods on at least six policy models, achieving an average performance that is 2% higher than the best-performing baseline. 2) Synthesizer-8B, demonstrates substantial improvements on larger models, such as performance gains of 9.6% for GLM-4-Plus and 1.6% for GPT-4o on GSM8k, even surpassing Llama3.1-70B-Instruct.
Synthesizer-8B generalizes well to unseen datasets, even when the training data does not include examples from specific benchmarks. For example, despite the absence of GSM8k and FeTaQA datasets in the training data, Synthesizer-8B achieves superior performance on these tasks, highlighting its effectiveness in handling unseen data.
Synthesizer-8B demonstrates robustness across structurally distinct models. Synthesizer-8B relies solely on Llama3-8B-Instruct to generate candidate responses and uses Llama3.1-70B-Instruct for generating replies during training. Despite this, our method still shows significant improvements across a wide range of policy models, including GLM-4-Plus, GPT-4o, Qwen2-7B-Instruct, and Qwen2.5-14B-Instruct, which are structurally and behaviorally distinct from the models used during training.
These findings highlight that our method effectively learns a generalized reasoning and synthesis strategy, enabling it to enhance the outputs of diverse policy models without overfitting to the specific characteristics of the models used for training data generation.
5.3 Ablation Studies
We conduct ablation studies to evaluate the impact of the process of CoT analysis of candidate responses and training on data generation pipeline.
Omitting CoT training affects robustness and generalization. In contrast to the original approach, where CoT plays a critical role, we directly utilize the Llama3.1-70B-Instruct model to provide answers to each question, generating high-quality responses. These responses are subsequently used as labels in the training data, thus skipping the CoT analysis of candidate responses. Table 2 indicates that skipping CoT training slightly improves average performance compared to the CoT-prompting baseline but remains less effective than Synthesizer-8B. Notably, on some policy models, such as Llama3.1-70B-Instruct and Qwen2.5-14B-Instruct, the performance of skipping CoT training is even worse than the CoT-prompting, while on Llama3-8B-Instruct testset it surpasses the Synthesizer-8B. This suggests that removing CoT can lead to overfitting to the candidate response generation model, thereby reducing transferability and generalization across different policy models.
Omitting training on generation data impacts reasoning and synthesis capabilities. To evaluate the effectiveness of our data generation pipeline, we directly utilized an untrained Llama3-8B-Instruct as the Synthesizer to regenerate candidate responses from various policy models on mathematical reasoning tasks. Table 2 illustrates the impact of removing synthetic data. On the MATH dataset, the untrained model shows some synthesis improvements compared to CoT-prompting for some certain policy models, such as performance gains of 8% for Llama3-8B-Instruct and 3.6% for Llama3.1-8B-Instruct. However, for stronger models such as Llama3.1-70B-Instruct and Qwen2.5-14B-Instruct, the performance through answer synthesis using the untrained model is even inferior to CoT-prompting. Similarly, on the GSM8k dataset, the untrained model failed to enhance most policy models, indicating that it still makes errors during reasoning and synthesis.
5.4 Scaling Experiments
We experiment with training data size and inference iterations to assess the scalability of our method.
Training Data Scaling.
Figure 3: Effects of training data size on the performance of Scalar RM and Ours across the MATH and GSM8k datasets. As the data size increases, the performance of Scalar RM initially improves but subsequently declines while our method exhibits a consistent increase.
We evaluate the impact of training data size on both Scalar RM and our method. Specifically, we train the mathematical reasoning tasks with varying amounts of training data: 2k, 4k, 8k, 16k, 32k, 64k, 128k, keeping all settings consistent with the main experiment. Figure 3 illustrates the average performance trends of these two methods on the mathematical tasks as the data size increases.
For our method, performance improves in a log-linear relationship with the data size and consistently outperforms Scalar RM across all settings. Due to the limitations of the available data, training is terminated at 128k data points.
In contrast, Scalar RM exhibits a different trend: its performance initially with increasing data size but subsequently declines. This decline may be attributed to the limited training data diversity, which is generated through multiple samplings of the original benchmark dataset, leading to a significant number of duplicate instructions. Since Scalar RM is trained using positive and negative pairs and produces scalar scores during inference, it is more prone to “overfitting” on these repeated instructions. Additionally, as each dataset is trained for the same epoch, this overfitting becomes more pronounced with larger datasets. In comparison, our generative approach is inherently more robust to such repetition, remaining unaffected by the lack of diversity in the training data.
Inference Scaling.
We further investigate the impact of varying the number of candidate responses on the performance of policy models, including Self-consistency, ArmoRM, Scalar RM and our method. For this analysis, we use Llama3-8B-Instruct and Qwen2-7B-Instruct as policy models, conducting 125 inferences on the MATH test set. To ensure consistency across experiments, 1, 5, and 25 samples are randomly selected from the candidate answers for comparison.
Since our method involves concatenating multiple candidate responses before inputting them into the model, input length constraints prevent the model from processing all candidates simultaneously when the number of candidate responses becomes too large. To address this, candidate answers are grouped into sets of five, with any remaining answers treated as a separate group. The synthesized outputs from each group are then further combined iteratively across groups.
Figure 4: Effects of inference scaling on the performance of Scalar RM and Ours across the MATH and GSM8k datasets.
As shown in the Figure 4, as the number of candidate responses increases, the performance of our method and Self-consistency improves steadily, indicating their robustness. In contrast, the performance of the ArmoRM and Scalar RM, both of which adopt Best-of-N methods, shows an initial improvement followed by a decline as the number of candidate answers grows. This decline can be attributed to the verifier being distracted by other solutions containing erroneous information while searching for the optimal solution, leading to a “reward hacking” phenomenon. This observation aligns with the description provided in Cobbe et al. (2021).
5.5 Analysis
We select Llama3-8B-Instruct as the policy model and conduct a comparative analysis of SC, ArmoRM, Scalar RM and our method on the MATH test set. For each sample, the policy model generates 5 candidate responses, and we compare the number of correct answers as well as the final results after applying the post-processing methods, as shown in Table 3. Notably, our method can synthesize a correct answer even when all candidate responses are incorrect. Figure 5 provides a detailed example.
Correct Count 0 1 2 3 4 5 Sum
SC 0 17 37 36 32 40 162
ArmoRM 0 36 35 22 31 40 164
Scalar RM 0 36 37 28 30 40 171
Synthesizer-8B 9 33 38 36 32 40 188
Table 3: Distribution of correct answers for different post-processing methods on Llama3-8B-Instruct model across 5 inferences.
6 Conclusion
This paper proposes CoT-based Synthesizer, a novel inference scaling strategy that integrates information from multiple candidate responses, enabling accurate answers. To enable a lightweight and cost-effective implementation, we introduce an automatic data generation pipeline, allowing smaller models to learn effectively from diverse data generated by large LLMs. Experiments across benchmarks demonstrate that CoT-based Synthesizer outperforms traditional methods like Best-of-N and Self-consistency.
Limitations
Grouping and Synthesis of Candidate Responses.
While our method supports the synthesis of multiple candidate answers, the input length limitation of the backbone model necessitates grouping the candidate answers for separate synthesis, followed by iterative synthesis of the grouped outputs. Although this approach leverages relationships among different candidate answers within the same group, it weakens the connections between candidates across different groups, and it also increases the overall inference cost. With the emergence of more advanced long-context LLMs, our method is expected to achieve more efficient and higher-quality solution generation by leveraging multiple candidate answers.
Inference Overhead.
Although our method processes multiple candidate answers simultaneously and performs synthesis in a single step, thereby significantly reducing the number of inference passes compared to the Best-of-N approach, which involves scoring and filtering each candidate answer, it still requires all candidate answers to be generated before synthesis can take place. This design introduces additional inference time overhead.
Appendix A License
Our research utilizes the training datasets from MATH and WikiTQ as foundational resources. These datasets are distributed under the Apache 2.0 license, which permits users to freely use, modify, reproduce, and share the software for both personal and commercial purposes.
In alignment with the principles of open access, we commit to publicly releasing the training data upon the acceptance of this work. The released data will be licensed under the CC BY-SA 4.0 license, ensuring its reuse and redistribution are permitted, provided that derivative works maintain the same licensing terms.
Appendix B Experiment Settings
B.1 Experimental Environment
Our experiments are conducted on a server running the Ubuntu 22.04 operating system, utilizing PyTorch version 2.4.0. The system is equipped with 8 NVIDIA A800 80GB GPUs, an Intel (R) Xeon (R) Platinum 8358 CPU, and 2048GB of RAM.
B.2 Baselines and Implementation
The policy models are divided into two types, including API-based models and open-source models. For the open-source models that require us to deploy them ourselves, we use PyTorch, Transformers, and vLLM Kwon et al. (2023) to load the models. For each policy model, we generate five candidate responses. The inference parameters can be referred to in Table  4. For detailed information about each policy model, please refer to Table 6.
Hyperparameters Value
Temperature 0.9
Top P 0.9
Max Tokens 1024
Frequency Penalty 0
Presence Penalty 0
Table 4: The hyperparameters of LLMs for candidate response generation.
For baseline methods, we first use the prompts specifically designed for each dataset to generate chain-of-thought (CoT) responses, with prompt details provided in Appendix C. The detailed information for the other baselines is introduced as follows:
  • Consistency-based Methods: For mathematical reasoning, we adopt the standard Self-consistency configuration. However, since TableQA involves open-ended responses, it is challenging to extract precise answers for majority voting, making it unsuitable for Self-consistency tasks. Therefore, we also utilize a variant of consistency-based methods, USC, across four datasets. For a fair comparison, we use Llama3.1-70B-Instruct as the base model for USC experiments, enabling majority voting among candidate answers under consistent experimental conditions.
  • Best-of-N Methods: We initially select ArmoRM-Llama3-8B-v0.1, which currently achieves the highest reasoning performance on the Reward Bench leaderboard based on Llama3-8B-Instruct. Additionally, we train a standard scalar reward model following the approach in  Cobbe et al. (2021). Specifically, for the MATH and WikiTQ datasets, to achieve a consistent training data size with our method, we use Llama3-8B-Instruct to generate 15 samples per data instance with a temperature of 0.9. Correct answers (including the gold answer) are treated as positive examples, while incorrect answers are treated as negative examples. This results in a dataset of 300k positive-negative pairs, which we use to train the scalar RM with Llama3-8B-Instruct.
  • Synthesis-Based Methods: LMCOR is a synthesis-based method designed for inference scaling. It generates candidate answers by synthesizing responses based on the gold answers in the dataset. We employ Llama3-8B-Instruct as the base model and Llama3.1-70B-Instruct as the policy model to generate candidate responses, keeping all other configurations consistent with the original method.
B.3 Training Implementation
Benchmark Train Size Extended Size Test Size
GSM8k - - 1319
MATH 12k 295k 500
FeTaQA - - 753
WikiTQ 18k 87k 633
Table 5: Benchmark dataset sizes for train and test.
Table 6: Information of baseline models.
For Scalar RM, LMCOR, and our method, we all use Llama-3-8B-Instruct as the base model.
For our method, we utilize Llama-3-8B-Instruct to generate five candidate responses for each sample. Subsequently, these candidates are subjected to CoT synthesis using Llama3.1-70B-Instruct. For the MATH training dataset, we perform an initial sampling of 50 iterations per sample. If none of these iterations yield a correctly synthesized answer, we apply LLM Repair and conduct a second round of 20 additional samplings. For the WikiTQ training dataset, we carry out an initial sampling of 20 iterations, followed by a second round of 10 additional samplings for samples that fail to produce a correctly synthesized answer in the first round. For the data filtering in the TableQA task, we use CritiqueLLM Ke et al. (2023), which is an LLM focused on evaluating the reasoning process and the consistency of answers with the benchmark set through scoring. We retained results with scores greater than or equal to 8 in order to obtain high-quality answers. Through this data synthesis pipeline, we have generated a comprehensive dataset consisting of 294k synthesized samples for the MATH dataset and 87k synthesized samples for the WikiTQ dataset. Detailed information is shown in Table 5. To ensure the reliability of our results, each experiment is conducted three times, and we report the average outcomes.
Following an extensive hyperparameter search over learning rates, we determine that an LR range of [5e-6, 2e-6, 5e-7] yields optimal performance for all training methods. Thus we choose LR = 2e-6 for all methods to obtain the best outcomes. We utilize the AdamW optimizer Loshchilov (2017) with decoupled weight decay regularization, setting the weight decay to 1e-2. Dropout is not applied during training.
In our experimental setup, we configure the per-GPU batch size to 1, with gradient accumulation over 16 steps, thereby achieving an effective batch size of 128 when distributed across 8 GPUs, each for 2 epochs. We set the maximum length of 4096 tokens and employ BF16 precision.
Appendix C Prompt for Inference
Prompt for Synthesizer inference [Instruction]
Please act as an excellent summarizer and summarize the following AI responses to the questions. Your summary should fully consider the connection between the question and AI responses, resulting in a correct, high-quality answer. In most cases, the same response that appears most often in the response may be the correct answer. If you find that there is no correct answer, please try to generate a correct answer yourself. Do not copy The candidate’s answer, give your summarized answer and reasons, and give the correct answer at the end of the sentence in the format: The answer is…
[The Start of Original Question]
{question}
[The End of Original Question]
[The Start of AI Responses]
{responses}
[The End of AI Responses]
Prompt for LLM Repair [Instruction] Please act as an excellent summarizer. Most of the candidate answers provided are wrong; you should find the right steps in these AI responses and try to combine these right steps to generate a new correct, high-quality answer. Give your summarized answer and reasons, and give the correct answer at the end of the sentence, in the format: The answer is… [The Start of Original Question]
{question}
[The End of Original Question]
[The Start of AI Responses]
{responses}
[The End of AI Responses]
Prompt for TableQA CoT-prompt Offer a thorough and accurate solution that directly addresses the Question outlined in the [Question]. Make sure that the last sentence is formatted as "the answer is XXX" and try to use words from the table or the question. ### [Table]: {table} ### [Question]: {question} ### [Solution]:
Prompt for MATH CoT-prompt As an expert problem solver, solve step by step the following mathematical questions. Make sure that the last sentence is formatted as "The answer is XXX" and the value of the correct answer is surrounded by boxed when giving your answer. [Problem]: {question} [Solution]:
Prompt for GSM8k CoT-prompt As an expert problem solver solve step by step the following mathematical questions. Make sure that the last sentence is formatted as "The answer is XXX" when giving your answer. Q: {question} A:
{mdframed}
[style=exampledefault,frametitle=An example of synthesizing a correct answer from all the wrong candidate responses] [Instruction]
Please act as an excellent summarizer. Below are some candidate answers generated by AI assistants. You need to do two tasks based on these responses: First, you should analyze these AI responses and find the right steps among them. Second, you should try to synthesize these right steps to generate a new correct, high-quality answer. If you can’t get a correct answer based on these answers, you should generate a correct answer. Give your summarized answer and reasons, and give the correct answer at the end of the sentence, in the format: The answer is…
[The Start of Original Question]
As an expert problem solver, solve step by step the following mathematical questions. Make sure that the last sentence is formatted as "The answer is XXX" and the value of the correct answer is surrounded by boxed when giving your answer.
### Problem:
What real values of $x$ are not in the domain of
$f(x)=\frac{1}{|x^{2}+3x-4|+|x^{2}+9x+20|}$?
### Solution:
[The End of Original Question]
[The Start of AI Responses]
Response 1:…… The answer is $\boxed{-2,-4}$.
[…]
Response 3:…… The answer is $\boxed{\{-4,-5\}}$.
Response 4:…… The answer is 1,4,5,41454- 1 , - 4 , - 5 , 4.
Response 5:…… The answer is $\boxed{-4}$.
[The End of AI Responses]
Response:
Based on the given AI responses, the correct answer can be determined by analyzing the connection between the question and the responses. The question asks for the real values of x that are not in the domain of the function f(x).
Response 1 and Response 5 provide the correct approach to solve the problem by setting the expression inside the absolute value symbols equal to zero and solving for x. However, Response 5 incorrectly states that x = 1 is a solution, which is not supported by the other responses.
Response 2 incorrectly states that the domain of f(x) is the set of all real numbers, which is not true since the function has specific values that make the expression inside the absolute value symbols equal to zero.
Response 3 and Response 4 provide similar solutions, but Response 4 incorrectly includes x = 4 and x = -1 as solutions, which are not supported by the other responses.
Considering the connection between the question and the responses, the correct answer is the values of x that make the expression inside the absolute value symbols equal to zero, which are x = -4 and x = -5 (from Response 3 and Response 4) and x = -2 and x = -4 (from Response 1). However, Response 3 and Response 4 do not agree on the value x = -5, and Response 1 and Response 5 do not agree on the value x = -2 and x = 1, respectively. Since Response 1, Response 3, and Response 5 provide a clear solution, the correct answer is the intersection of the values provided by these responses.
The answer is $\boxed{-4}$.
Figure 5: An example of synthesizing a correct answer from all the wrong candidate responses generated by Llama3-8B-Instruct on MATH. The ellipsis indicates responses to the MATH instruction by CoT-prompting.
{mdframed}
[style=exampledefault,frametitle=An example of synthesizing a correct answer from all the wrong candidate responses] [Instruction]
Please act as an excellent summarizer. Below are some candidate answers generated by AI assistants. You need to do two tasks based on these responses: First, you should analyze these AI responses and find the right steps among them. Second, you should try to synthesize these right steps to generate a new, correct, high-quality answer. If you can’t get a correct answer based on these answers, you should generate a correct answer. Give your summarized answer and reasons, and give the correct answer at the end of the sentence in the format: The answer is…
[The Start of Original Question]
Offer a thorough and accurate solution that directly addresses the Question outlined in the [Question]. Make sure that the last sentence is formatted as "the answer is XXX" and try to use words from the table or the question.
### [Table Description]
The table consists of match records for a certain event, featuring columns for Match ID, Date, Venue, Opponents, and Score. The Match ID includes a code representing the stage of the tournament and a sequential number. The Date column indicates the year of the match, and the Venue column contains the location of the match. The Opponents column lists the competing teams, while the Score column captures the outcome of each match.
### [Table]
Match,Date,Venue,Opponents,Score
GL-A-1,2006..,[[]],[[]],-
GL-A-2,2006..,[[]],[[]],-
GL-A-3,2006..,[[]],[[]],-
GL-A-4,2006..,[[]],[[]],-
GL-A-5,2006..,[[]],[[]],-
GL-A-6,2006..,[[]],[[]],-
Quarterfinals-1,2006..,[[]],[[]],-
Quarterfinals-2,2006..,[[]],[[]],-
Semifinals-1,2006..,[[]],[[]],-
Semifinals-2,2006..,[[]],[[]],-
### [Question]
How many matches are from 2006?
[The End of Original Question]
[The Start of AI Responses]
Response 1: To determine the number of matches from 2006, we need to examine the ’Date’ column in the given table. After reviewing the table, we find that all the matches listed have a date in 2006. Specifically, there are 12 matches with dates in 2006, as indicated by the ’GL-A’ matches and the tournament stages listed in the ’Match’ column. Therefore, the answer is 12 matches are from 2006.
Response 2: To find the number of matches from 2006, we need to count the rows where the Date column contains the year 2006. In the provided table, there are 10 rows, and 8 of those rows have the Date column indicating the year 2006. Therefore, the answer is 8 matches are from 2006.
Response 3: ……The answer is 8.
Response 4:……Therefore, the answer is 8 matches are from 2006.
Response 5:……The number of matches from 2006 is 8.
[The End of AI Responses]
Response:
After reviewing the AI responses, it is clear that the majority of the responses indicate that there are 8 matches from 2006. However, upon closer inspection of the provided table, all rows listed have a date in 2006, which totals 10 matches (including the 8 matches from the "GL-A" series and the 2 quarterfinal matches). The discrepancy in the AI responses might be due to an oversight in counting the quarterfinal matches.
The answer is 10 matches are from 2006.
Figure 6: An example of synthesizing a correct answer from all the wrong candidate responses generated by Qwen2-7B-Instruct on WikiTQ. The ellipsis indicates responses to the WikiTQ instruction by CoT-prompting.